OpenStudy (pi123):

Help please!!! Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y).

OpenStudy (pi123):

i came up with this but im not sure its right : 20y+x=50

OpenStudy (pi123):

@ganeshie8 can you help me?

OpenStudy (anonymous):

y = 20 + x x + y = 50 They said write "a pair" of linear equations, so you will need 2 equations

OpenStudy (pi123):

Yeah i realized that now thank you lots! do you think you can help me with one more? @Purplerainbowcherry

OpenStudy (anonymous):

sure! :)

OpenStudy (pi123):

thank you! :) let me go grab the question one sec. oh and i just fanned you.

OpenStudy (pi123):

A system of linear inequalities is shown below: x + 2y < -5 x - y > 0 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

hahaha no prob, thanks :)

OpenStudy (pi123):

i know that on the graph it would be a dotted line and for the first equation you have to shade below, and for the second equation you have to shade above. But I dont know how to really describe how to graph it.

OpenStudy (anonymous):

so what exactly arent you able to understand?

OpenStudy (pi123):

i dont know how to really explain how to graph it. i struggle with graphing inequalities. sorry i just realized you replied.

OpenStudy (pi123):

could you still help me? @Purplerainbowcherry

OpenStudy (anonymous):

you have to find points to be able to draw the 2 lines, make a table for each line and decide a set of values for x and find their corresponding y values

OpenStudy (pi123):

thank you!!!